OpenStudy (anonymous):

Write an equation of a line parallel to y = 1 over 4x − 6 in slope-intercept form that passes through the point (−1, 5). y= 1 over 4x − 1 <- my answer correct me if I'm wrong please y= 1 over 4x + 5.25 y= −4x + 5 y= −4x − 1.25

OpenStudy (anonymous):

use this.. (y-y1)=(x-x1) herre m=slope

OpenStudy (nurali):

if y = 1 over 4x − 6, m = (1/4) Using point-slope form, (y-y1)=m(x-x1) P(-1,5) y -5 = (1/4)(x + 1) y = (1/4)x + 1/4 + 20/4 y = (1/4)x + 21/4